# A Gift from the Integration of Discriminative and Diffusion-based Generative Learning: Boundary Refinement Remote Sensing Semantic Segmentation

### v1

@@ IV-C3 Evaluation on Cross-Phase Building Segmentation
TABLE IV: Quantitative results (%) in the post-disaster phase of the WHU Building dataset.
Method DiscModel Backbone Building Kappa OA WFm (3px B)
IoU F1score
DDP ConvNext-t 93.03 96.39 95.75 98.91 69.44
DeepLabV3+ ResNet-50 93.00 96.37 95.73 98.91 68.41
@@
TABLE V: Quantitative results (%) in the post-reconstruction phase of the WHU Building dataset.
Method DiscModel Backbone Building Kappa OA WFm (3px B)
IoU F1score
DDP ConvNext-t 92.97 96.35 95.52 98.64 72.23
DeepLabV3+ ResNet-50 91.70 95.67 94.69 98.40 66.59
@@
In the initial post-disaster phase (Table IV), IDGBR brought significant improvements across all baseline models in every evaluation metric. For DeepLabV3+, SegFormer, and DINOv2, the WFm increased by 6.55%, 4.75%, and 8.95%, respectively. IDGBR maintained stable performance in cross-phase testing (Table V). Taking DeepLabV3+ as an example, its WFm improved from 68.41% to 74.96% (+6.55%) and IoU rose from 93.00% to 94.37% in the post-disaster phase; while maintaining 71.87% WFm (+5.28%) and achieving 92.73% IoU (from 91.70% baseline) in the post-reconstruction phase. At the macro level, IDGBR yielded average improvements of 5.21% in WFm and 1.21% in IoU during the post-disaster phase, while achieving gains of 4.20% (WFm) and 0.95% (IoU) in the post-reconstruction phase, which demonstrates IDGBR’s robust cross-temporal adaptability.
@@ IV-D Evaluation on Multi-Class Segmentation Tasks
In multi-class segmentation, we employ Potsdam and Vaihingen as benchmark datasets, conducting comparative analyses of model performance in complex boundary scenarios (such as building edges, fragmented vegetation boundaries, vehicle contours, etc.). This approach validates the effectiveness of our method in optimizing boundary regions within multi-class segmentation tasks.
IV-D1 Evaluation on Potsdam Dataset
Figure 7: Qualitative results of different methods on multi-class segmentation dataset.
TABLE VI: Quantitative results (%) of different methods on the Potsdam dataset.
Method DiscModel- Backbone Impervious surfaces Building Low vegetation Tree Car mean F1 mIoU WFm (3px B)
DDP ConvNext-t 93.13 96.67 86.56 86.24 91.3 90.78 83.36 41.11
DeepLabV3+ ResNet-50 93.18 96.95 86.51 86.07 91.63 90.87 83.53 37.40
@@ IV-D1 Evaluation on Potsdam Dataset
From the experimental results in Table VI, IDGBR framework achieves superior boundary refinement by synergistically integrating the semantic discriminability of discriminative models with the edge delineation strengths of diffusion models, outperforming both conventional discriminative baselines and DDP in segmentation precision. Notably, IDGBR achieved substantial improvements in WFm metrics across all tested models, with gains ranging from 2.88% to 8.12%, where DINOv2 demonstrated the most remarkable enhancement from 35.12% to 43.24%. Meanwhile, the method maintained stable mIoU performance across different architectures, demonstrating that IDGBR successfully enhances boundary precision without compromising overall segmentation quality.
IV-D2 Evaluation on Vaihingen Dataset
TABLE VII: Quantitative results (%) of different methods on the Vaihingen dataset
Method DiscModel- Backbone Impervious surfaces Building Low vegetation Tree Car mean F1 mIoU WFm (3px B)
DDP ConvNext-t 89.50 94.53 81.68 88.23 81.93 87.17 77.60 45.68
DeepLabV3+ ResNet-50 88.92 94.04 80.27 87.45 78.98 85.93 75.76 40.38
@@ IV-D2 Evaluation on Vaihingen Dataset
The experimental results on the Vaihingen dataset, as shown in Table VII, demonstrate that compared to discriminative baseline models, our framework effectively improves both WFm and mIoU across different model architectures. Regarding the WFm metric, our framework yielded improvements of 4.76%, 1.28%, 1.28%, and 2.40% for DeepLabV3+, SegFormer, DINOv2, and LSK, respectively, further validating our method’s enhancement of boundary representation capabilities in multi-class remote sensing segmentation tasks.
@@ V Discussion
V-A Enhancing Semantic Awareness in Diffusion Inference: Effects of Conditional Guidance and Cubic Timestep Sampling
TABLE VIII: Qualitative results (%) of our framework under different guidance conditions on CHN6-CUG and Potsdam datasets.
Setting Requirements CHN6-CUG Potsdam
@@
In the proposed IDGBR framework, we use remote sensing images and rough labels as conditional guidance, and introduce a cubic timestep sampling strategy to enhance semantic representation capabilities. To systematically assess the contribution of each component, we designed three ablation experiments, which can be divided into three groups: (1). Experiment A: using only remote sensing images as guidance; (2). Experiment B: using remote sensing images and rough labels jointly guided by the proposed Conditional Guidance Module; (3). Experiment C: introducing cubic timestep sampling technique into the timestep sampling based on Experiment B. As shown in Table VIII, we compared three different strategies: “Image” refers to the use of remote sensing image guidance, “Rough” refers to the use of rough label guidance, and “Cubic” refers to the additional use of the cubic timestep sampling technique; check marks and crosses are used to indicate whether the corresponding feature is enabled.
@@
By comparing the three sets of experimental results, Experiment A recorded the lowest WFm values (33.66% on CHN6-CUG; 30.39% on Potsdam) along with significantly reduced mF1 and mIoU versus other groups. This indicates that without rough label guidance, the model struggles to capture object boundaries. In Experiment B, WFm of the two datasets increased to 37.64% and 40.34% respectively, and the mIoU also improved substantially. This demonstrates that rough labels as additional semantic guidance play a crucial role in boundary modeling. In Experiment C, WFm in the Potsdam dataset further increased to 42.59%, with other metrics also showing an upward trend. This indicates that cubic timestep sampling strengthens feature discrimination at semantic boundaries across multiple classes. However, in the binary segmentation CHN6-CUG dataset, this strategy led to a slight decrease in all metrics. This phenomenon can be attributed to differences in task requirements. Binary segmentation requires precisely capture significant feature differences between local specific targets and their backgrounds, without considering complex feature relationships among other irrelevant targets in the global scene. The cubic sampling strategy, by increasing sampling frequency in high-noise regions, forces the diffusion model to allocate more attention to subtle feature differences among various potential targets throughout the entire scene, significantly enhancing the model’s discriminative ability in multi-class segmentation tasks. However, this shift in feature learning direction interferes with binary segmentation task learning, causing the model to no longer focus on learning the most significant discriminative features between a single target and its background, thereby weakening the model’s precision in single-target discrimination. Experiments validate that integrating the Conditional Guidance Module with cubic timestep sampling in IDGBR significantly boosts boundary quality and the processing of complex semantic structures in multi-class segmentation tasks.
V-B Discussion on Applicability and Limitations
We categorize errors in rough labels into two types: semantic boundary errors and semantic content errors. For semantic boundary errors in rough labels, experimental results demonstrate the outstanding capability of our proposed method to optimize semantic boundaries in both binary and multi-class classification tasks. For semantic content errors in rough labels, our method exhibits dependence on the completeness of information in correct patches.
Based on the completeness of information in correct patches, this dependency can be divided into three progressive levels:
@@ V-B Discussion on Applicability and Limitations
When semantic patches exhibit partial segmentation omissions or commissions, yet the correct patch regions can still approximately represent or deduce complete information through semantic adjacency relationships, our method effectively repairs erroneous semantics. Fig. 8 showcases complete semantic patch restoration cases. (i) In (a-2), the rough label shows massive segmentation omissions on the left side of the road (ROI), but the scattered correct patches are sufficient to reveal the length and width of the road area. The spatial distribution patterns of roads have already been learned by the generative model during its training on the road annotation dataset. During inference, these scattered patches are sufficient to activate the road geometric structure prior in the model weights, thereby completing the global supplementation of the semantic patch. (ii) In (b-2), the spatial position of the road patch is basically correct, but its internal structure appears fragmentation. The road continuity patterns learned by the model during training effectively correct the local errors within the patch. (iii) In (c-2), the spatial positions of impervious surface and low vegetation patches within the ROI region are generally correct, yet boundary delineations between patches show imprecision. The model addresses these edge-related semantic inaccuracies through boundary refinement mechanisms. (iv) In (d-2), the impervious surface patch in the ROI region was misclassified as building categories, but the model corrected these scattered classification errors by utilizing the continuity patterns of impervious surfaces.
@@
When semantic patches exhibit partial segmentation omissions or commissions, yet correct patch regions can only infer localized information through semantic adjacency relations, our method only correct the relevant local semantics. Fig. 9 showcases partial semantic patch restoration. (i) In (a-2), the majority of the non-cultivated land within the ROI were misclassified. Although scattered background patches activated the farmland boundary continuity patterns learned by the model from the farmland dataset, this was insufficient to support complete reconstruction of the patch width. Consequently, the optimization could only refine the slender structure of the non-cultivated land. (ii) In (b-2), there are only a few road semantic patches at the bottom of the ROI. The road geometric patterns learned by the model during training can geometrically correct this area but are insufficient to infer the length information of the road. (iii) In (c-2), building patches in the ROI were misclassified as background. The building area in the image was under construction, with texture features significantly different from surrounding buildings, making it difficult for the model to identify incorrect patches by comparing the image with rough labels. (iv) In (d-2), low-rise buildings on the left side of the ROI region were misclassified as impervious surfaces due to radiometric distortion caused by shadow interference from adjacent high-rise structures on the right. The model could correct the misclassified patches on the right through image and coarse label comparison, but the shadow-affected areas on the left side remained uncorrected.
Figure 10: Examples of irreparable incorrect semantic patches.
@@
When semantic patches exhibit complete omissions or commission errors, and the semantic content cannot be reliably inferred from adjacent patches, our method is unable to rectify these inaccuracies. Fig. 10 showcases instances where incorrect semantic patches cannot be restored. (i) In (a-2) , correctly classified semantic patches within the ROI are fragmented and lack farmland characteristics and spatial continuity, resulting in their removal as noise by the model. (ii) In (b-2), the gap between the left and right buildings in the ROI with ambiguous boundaries, leading the model to misidentify them as a single connected region despite comparing the image with rough labels. (iii) In (c-2), impervious surface patches in the ROI were misclassified as background. Despite comparing the image with coarse labels, the model corrected only the geometric errors while failing to address the semantic misclassification. (iv) In (d-2), building patches in the ROI were misclassified as impervious surfaces. The area shows open-air renovated balconies that closely resemble impervious surfaces in texture and color, causing the model to fail to detect the semantic errors in the rough labels.
V-C Speeding Up Representation Learning through Incipient Alignment Strategies

### v2

@@ IV-C3 Evaluation on Cross-Phase Building Segmentation
TABLE IV: Quantitative comparison (%) of our proposed method against various general segmentation methods on the post-disaster phase of the WHU Building dataset.
Method DiscModel Backbone Building Acc WFm (3px B)
IoU F1-score
DeepLabV3+ ResNet-50 93.00 96.37 98.91 68.41
@@
TABLE V: Quantitative comparison (%) of our proposed method against various general segmentation methods on the post-reconstruction phase of the WHU Building dataset.
Method DiscModel Backbone Building Acc WFm (3px B)
IoU F1-score
DeepLabV3+ ResNet-50 91.70 95.67 98.40 66.59
@@
In the initial post-disaster phase (Table IV), IDGBR yielded significant improvements across all baseline models in every evaluation metric. For DeepLabV3+, SegFormer, and DINOv2, the WFm increased by 6.55%, 4.75%, and 8.95%, respectively. IDGBR maintained stable performance in cross-phase testing (Table V). Taking DeepLabV3+ as an example, its WFm improved from 68.41% to 74.96% (+6.55%) and IoU rose from 93.00% to 94.37% in the post-disaster phase; while maintaining 71.87% WFm (+5.28%) and achieving 92.73% IoU (from 91.70% baseline) in the post-reconstruction phase. In terms of overall performance, IDGBR yielded average improvements of 5.21% in WFm and 1.21% in IoU during the post-disaster phase, while achieving gains of 4.20% (WFm) and 0.95% (IoU) in the post-reconstruction phase, which demonstrates IDGBR’s robust cross-temporal adaptability.
@@ IV-D Evaluation on Multi-Class Segmentation Tasks
In multi-class segmentation tasks, we employ Potsdam and Vaihingen as benchmark datasets to conduct a comparative analysis of model performance in complex boundary scenarios.
As presented in Table VI, on the Potsdam dataset, the IDGBR framework demonstrates superior boundary refinement by synergistically integrating the semantic discriminability of baseline models with the edge delineation capabilities of diffusion models. Quantitatively, IDGBR yields significant WFm improvements across all tested architectures, with gains ranging from 2.88% to 8.12%. Notably, the DINOv2 baseline exhibits the most remarkable enhancement, where our method elevates its WFm score from 35.12% to 43.24%.
Similarly, as shown in Table VII, results on Vaihingen further confirm the effectiveness of the proposed method. Compared with baseline models, our framework effectively improves the WFm metric across different model architectures. For instance, IDGBR yielded WFm improvements of 4.76%, 1.28%, 1.28%, and 2.40% for DeepLabV3+, SegFormer, DINOv2, and LSK, respectively. These results indicate that IDGBR significantly enhances boundary representation without compromising overall segmentation quality.
TABLE VI: Quantitative comparison (%) of our proposed method against various general segmentation methods on the Potsdam dataset.
Method DiscModel Backbone mean F1 mIoU WFm (3px B)
DeepLabV3+ ResNet-50 90.87 83.53 37.40
@@
TABLE VII: Quantitative comparison (%) of our proposed method against various general segmentation methods on the Vaihingen dataset.
Method DiscModel Backbone mean F1 mIoU WFm (3px B)
DeepLabV3+ ResNet-50 85.93 75.76 40.38
@@ V Discussion
V-A Comparative Analysis with Specialized Boundary Optimization Methods
To rigorously benchmark the boundary refinement capability of IDGBR, we conducted a comprehensive comparison against representative boundary optimization methods. Specifically, comparisons are made against two categories: (1). Discriminative boundary optimization methods, including general vision models (PointRend [5], Mask2Former [37]) and the remote-sensing-specific MSEONet [24]; and (2). Diffusion-inspired discriminative segmentation methods, represented by DDP [43]. Under this framework, IDGBR is deployed to refine the rough labels generated by SegFormer. Extensive experiments across the aforementioned three binary and two multi-class datasets substantiate that IDGBR yields significant performance margins, even against these methods explicitly engineered for boundary optimization.
TABLE VIII: Quantitative results (%) of methods specialized for boundary optimization on the CHN6-CUG road dataset.
Method Road Acc WFm (3px B)
IoU F1-score
PointRend 62.79 77.11 97.45 35.36
Mask2Former 63.88 77.96 97.52 37.61
MSEONet 61.69 76.31 97.41 31.92
DDP 64.19 78.19 97.59 36.52
IDGBR (ours) 64.54 78.45 97.50 37.64
TABLE IX: Quantitative results (%) of methods specialized for boundary optimization on the FGFD.
Method Farmland Acc WFm (3px B)
IoU F1-score
PointRend 81.86 92.84 91.66 38.76
Mask2Former 85.94 92.44 91.27 44.27
MSEONet 87.06 93.08 91.93 35.13
DDP 87.44 93.30 92.16 42.40
IDGBR (ours) 87.66 93.42 92.39 49.78
TABLE X: Quantitative results (%) of methods specialized for boundary optimization on the post-disaster phase of the WHU Building dataset.
Method Building Acc WFm (3px B)
IoU F1-score
PointRend 92.95 96.34 98.90 68.44
Mask2Former 93.51 96.65 98.99 71.33
MSEONet 90.36 94.94 98.48 58.34
DDP 93.03 96.39 98.91 69.44
IDGBR (ours) 94.72 97.29 99.18 75.91
TABLE XI: Quantitative results (%) of methods specialized for boundary optimization on the post-reconstruction phase of the WHU Building dataset.
Method Building Acc WFm (3px B)
IoU F1-score
PointRend 91.16 95.37 98.28 66.99
Mask2Former 92.16 95.92 98.49 70.98
MSEONet 88.34 93.81 97.72 56.67
DDP 92.97 96.35 98.64 72.23
IDGBR (ours) 93.06 96.41 98.67 73.95
In binary segmentation scenarios, IDGBR consistently demonstrates superior boundary refinement across datasets with diverse characteristics. As evidenced in Table VIII, it effectively outperforms boundary optimization methods on the CHN6-CUG road dataset. This advantage is further amplified on the FGFD dataset (Table IX), where severe spectral ambiguity exists along boundaries. Here, IDGBR achieves the best IoU (87.66%) and F1-score (93.42%), while its WFm (49.78%) substantially leads all competitors, demonstrating the robustness of our method in boundary optimization. Furthermore, IDGBR exhibits strong cross‑temporal generalization on the WHU Building dataset (Tables X and XI). In particular, in the challenging post‑disaster phase, it attains a WFm of 75.91%, surpassing the best discriminative baseline Mask2Former by over 4.5%.
TABLE XII: Quantitative results (%) of methods specialized for boundary optimization on the Potsdam dataset.
Method mean F1 mIoU WFm (3px B)
PointRend 90.56 83.00 35.87
Mask2Former 91.26 84.21 40.79
MSEONet 90.05 82.17 34.28
DDP 90.78 83.36 41.11
IDGBR (ours) 90.78 83.36 42.59
TABLE XIII: Quantitative results (%) of methods specialized for boundary optimization on the Vaihingen dataset.
Method mean F1 mIoU WFm (3px B)
PointRend 86.57 76.68 41.74
Mask2Former 86.72 76.88 43.23
MSEONet 86.12 76.04 39.71
DDP 86.42 76.51 45.05
IDGBR (ours) 87.52 78.17 45.04
In multi-class segmentation scenarios, IDGBR maintains a significant advantage in boundary optimization. As presented in Tables XII and XIII, on the Vaihingen dataset, our method achieves a WFm of 45.04%, comparable to the top-performing DDP. Meanwhile, on the Potsdam dataset, it attains a leading WFm of 42.59%.
V-B Enhancing Semantic Awareness in Diffusion Inference: Effects of Conditional Guidance and Cubic Timestep Sampling
TABLE XIV: Quantitative results (%) of our framework under different guidance conditions on CHN6-CUG and Potsdam datasets.
Requirements CHN6-CUG Potsdam
@@ V-A Comparative Analysis with Specialized Boundary Optimization Methods
In the proposed IDGBR framework, we use remote sensing images and rough labels as conditional guidance, and introduce a cubic timestep sampling strategy to enhance semantic representation capabilities. To systematically assess the contribution of each component, we designed three ablation experiments, which can be divided into three groups: (1). Experiment A: using only remote sensing images as guidance; (2). Experiment B: using remote sensing images and rough labels jointly guided by the proposed Conditional Guidance Module; (3). Experiment C: introducing cubic timestep sampling technique into the timestep sampling based on Experiment B. As shown in Table XIV, we compared three different strategies: “Image” refers to the use of remote sensing image guidance, “Rough” refers to the use of rough label guidance, and “Cubic” refers to the additional use of the cubic timestep sampling technique; check marks and crosses are used to indicate whether the corresponding feature is enabled.
@@
By comparing the three sets of experimental results, Experiment A recorded the lowest WFm values (33.66% on CHN6-CUG; 30.39% on Potsdam) along with significantly reduced mF1 and mIoU versus other groups. This indicates that without rough label guidance, the model struggles to capture object boundaries. In Experiment B, WFm of the two datasets increased to 37.64% and 40.34% respectively, and the mIoU also improved substantially. This demonstrates that rough labels as additional semantic guidance play a crucial role in boundary modeling. In Experiment C, WFm in the Potsdam dataset further increased to 42.59%, with other metrics also showing an upward trend. This indicates that cubic timestep sampling strengthens feature discrimination at semantic boundaries across multiple classes. However, on the CHN6-CUG binary segmentation dataset, this strategy resulted in a slight decline across all evaluation metrics. Our extended analysis in the Appendix indicates that the cubic sampling strategy is best viewed as a condition-sensitive, optional optimization technique, particularly for binary segmentation.
V-C Discussion on Applicability and Limitations
To quantify the model’s repair efficacy on rough predictions, we constructed a partition-based evaluation metric derived from the discrepancy between the rough prediction $M_{r}$ and the ground truth $M_{t}$. First, we partition the rough prediction space into the mask of True predicted Region $M_{RTm}$ and the mask of False predicted Region $M_{RFm}$:
$M_{RTm}$ $=M_{t}\cap M_{r}$ (18)
$M_{RFm}$ $=(M_{t}\cup M_{r})-(M_{t}\cap M_{r})$
Based on this partition, we calculate the IoU of the refined prediction $M_{re}$ within these two specific regions:
$IoU_{RTm}$ $=\frac{|M_{t}\cap M_{re}\cap M_{RTm}|}{|(M_{t}\cap M_{RTm})\cup(M_{re}\cap M_{RTm})|}$ (19)
$IoU_{RFm}$ $=\frac{|M_{t}\cap M_{re}\cap M_{RFm}|}{|(M_{t}\cap M_{RFm})\cup(M_{re}\cap M_{RFm})|}$
where $IoU_{RTm}$ measures the model’s capability to preserve true predicted regions; an excessively low value indicates significant degradation of correct regions. $IoU_{RFm}$ evaluates the model’s ability to rectify false predicted regions, where a positive value indicates effective repair. Integrating both preservation and rectification capabilities, we define the final improvement score $IoU_{imp}$ by introducing a penalty term $k$ to balance their weights:
$IoU_{imp}=IoU_{RFm}-k\times(1-IoU_{RTm}),\quad k=1$ (20)
Based on the value of $IoU_{imp}$, we categorize repair efficacy into three levels:
  • Error Repair ($IoU_{imp}<0$): Implies that the damage to correct regions outweighs the repair benefits.
  • Partial Repair ($0\leq IoU_{imp}<80\%$): Indicates effective repair of some errors without causing severe damage to correct features.
  • Complete Repair ($IoU_{imp}\geq 80\%$): Represents efficient correction of erroneous regions while perfectly preserving the features of correct regions.
We recalculated the quantitative results for the three study scenarios using the $IoU_{imp}$ metric, applying it specifically to the Regions of Interest (ROI) in the diagrams. The following discussion analyzes and interprets these refined results as presented in the figure.
@@ V-C Discussion on Applicability and Limitations
When semantic patches exhibit partial segmentation omissions or commissions, yet the correct patch regions can still approximately represent or deduce complete information through semantic adjacency relationships, our method effectively repairs erroneous semantics. Fig. 7 showcases complete semantic patch restoration cases. (i) In (a-2), the refined result rectifies commission errors in the rough label by drawing on learned structure features, achieving an $IoU_{imp}$ of 85.02% for the ROI. (ii) In (b-2), the spatial position of the road patch is basically correct, but its internal structure appears fragmented. The road continuity patterns learned by the model during training effectively correct the local errors within the patch. (iii) In (c-2), the rough label misclassified impervious surfaces within the ROI as buildings. However, this region exhibits high feature homogeneity and spatial continuity with adjacent correctly labeled impervious surfaces. Leveraging these contextual consistency cues, the refined result effectively rectified the semantic error and regularized the boundaries, achieving a repair efficacy of $IoU_{imp}=92.37\%$. (iv) In (d-2), the impervious surface patch in the ROI region was misclassified as building categories, but the model corrected these scattered classification errors by utilizing the continuity patterns of impervious surfaces.
@@
When semantic patches exhibit partial segmentation omissions or commissions, yet correct patch regions can only infer localized information through semantic adjacency relations, our method only corrects the relevant local semantics. Fig. 8 showcases partial semantic patch restoration. (i) In (a-2), the rough label exhibits massive segmentation omissions in the ROI road area. Although learned geometric continuity priors successfully connected fragmented patches longitudinally into a complete linear structure, the scarcity of lateral information constrained accurate width inference. Consequently, the optimization achieved topological connectivity repair but failed to fully restore the original width, resulting in a partial repair efficacy of $IoU_{imp}=47.95\%$. (ii) In (b-2), there are only a few road semantic patches at the bottom of the ROI. The road geometric patterns learned by the model during training can geometrically correct this area but are insufficient to infer the length information of the road. (iii) In (c-2), the impervious surface in the ROI was misclassified as a building. Its high spectral similarity to adjacent buildings prevented semantic correction. Therefore, the model was limited to local geometric refinement. (iv) In (d-2), low-rise buildings on the left side of the ROI region were misclassified as impervious surfaces due to radiometric distortion caused by shadow interference from adjacent high-rise structures on the right. The model could correct the misclassified patches on the right through image and coarse label comparison, but the shadow-affected areas on the left side remained uncorrected.
Figure 9: Examples of failed semantic patch restoration.
@@
When semantic patches exhibit complete omissions or commission errors, and the semantic content cannot be reliably inferred from adjacent patches, our method is unable to rectify these inaccuracies. Fig. 9 showcases instances where incorrect semantic patches cannot be restored. (i) In (a-2), the majority of non-cultivated land within the ROI was misclassified. Although the model repaired topological interruptions based on continuity priors, the lack of localization cues caused the reconstructed trajectory to diverge from the true background and encroach into farmland. This significant spatial deviation resulted in negative optimization, with a final score of $IoU_{imp}=-11.44\%$. (ii) In (b-2), a white vehicle was misclassified due to its spectral similarity to buildings, which made semantic correction very difficult. The model was therefore limited to erroneously refining the contours of the spurious artifact. (iii) In (c-2), impervious surface patches in the ROI were misclassified as background. Despite comparing the image with rough label, the model corrected only the geometric errors while failing to address the semantic misclassification. (iv) In (d-2), building patches in the ROI were misclassified as impervious surfaces. The area shows open-air renovated balconies that closely resemble impervious surfaces in texture and color, causing the model to fail to detect the semantic errors in the rough label, resulting in negative optimization with $IoU_{imp}=-16.05\%$.
V-D Speeding Up Representation Learning through Incipient Alignment Strategies